40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 5 13:55:08 2011, kiwamu, Update, LSC, MI locking : calibration of BS and ITMs actuators MICH_OL.png
    Reply  Thu May 5 14:40:14 2011, Koji, Update, LSC, MI locking : calibration of BS and ITMs actuators 
       Reply  Thu May 5 15:28:23 2011, kiwamu, Update, LSC, Re: MI locking : calibration of BS and ITMs actuators 
Message ID: 4638     Entry time: Thu May 5 13:55:08 2011     Reply to this: 4639
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: MI locking : calibration of BS and ITMs actuators 

EDIT by KI on 15/5/2011:

The calibration of MICH error signal was wrong by a factor of 2.

 

The open loop transfer functions of the Michelson locking have been measured.

The purpose of this excise is to calibrate the coil-magnet actuators on BS and ITMs.

The estimated actuation coefficients are :

 BS = 3.69e-08 [m/counts]
 ITMX = 8.89e-09 [m/counts]
 ITMY = 9.22e-09 [m/counts] 

I guess the accuracy is something like 5 % because the calibration of the MI optical gain relies on a peak-to-peak measurement.

A next step is to calibrate the PRM actuator and the PRC optical gain.


 

(measurement)

The Michelson was locked with different actuators in every measurement. I locked the Michelson to the dark fringe with BS, ITMX and ITMY in each time.

The measured peak-to-peak value in the error signal was 20.2 counts, corresponding to a sensor gain of 5.96e+07 [counts/m]. Note that I used AS55_Q for the locking.

After locking the MI I took the open loop transfer function by injecting broadband noise from DTT.

Then the data were fitted coarsely.  In the fitting I used the resonant frequencies that Leo reported recently (http://blue.ligo-wa.caltech.edu:8000/40m/Mechanical_Resonances).

The Q-values are assumed to be 5 because of the local dampings. As a result the fitting gives us the actuator coefficients.

Here is a plot showing the measured open loop transfer functions. The solid lines represent the fittings.

MICH_OL.png

 

(by the way)

- The delay time including ADC/DAC and RFM looks quite big. According to the fitting the delay is something like 600 usec.

This is about two times larger than the one reported before (see this entry). I will re-measure it with empty filters.

ELOG V3.1.3-